OpenStudy (anonymous):

Can someone please explain how to solve this type of question? Determine the maximum and minimum value of the function f(x)=3x3^x - 1 (I will write out equation)

OpenStudy (anonymous):

\[f(x)=3x3^x-1\]

OpenStudy (anonymous):

i got f'(x)=3(3^x)+3x(ln3)^3^x as my derivative but I do not know how to simplify it

OpenStudy (anonymous):

You're on the right track by finding the derivative, which should be \[3*3^x + 3x(3^xln3)\] Now what you need to do is set this derivative equal to 0, and that will give you the x value.

OpenStudy (anonymous):

would the x value just be zero?

OpenStudy (anonymous):

Try plugging it in and see. If you plug in 0, then you get 3. What we're looking for is when the derivative is equal to 0.

OpenStudy (anonymous):

oh ok so order of operations? I think i'm doing this wrong so far: \[0=3(3^x)+3x(3xln3)\] \[0=9^x+9x^23xln3\]

OpenStudy (anonymous):

i dont know how to isolate x

OpenStudy (anonymous):

Your derivative isn't quite right, it should be \[3\cdot3^x+3x\cdot3^xln3\] Maybe that'll help.

OpenStudy (anonymous):

Hint: You have a common factor of 3*3^x in both terms. Factor that out

OpenStudy (anonymous):

Oh thanks. So now I have this as my x value: -1/ln3

OpenStudy (anonymous):

would i now find the second derivative and then substitute the x value into the function of the second derivative?

OpenStudy (alekos):

No. that's the end of the problem. This is the minimum value of the function. There is no maximum.

OpenStudy (anonymous):

ok thanks! also, how would you know if there is no minimum or no maximum?

OpenStudy (alekos):

just plot the graph using wolfram